# Unifying area and unit-level small area estimation through calibration

### v1

@@ 2 Area-level FH model
[…]
@@ 6 MSE of PEBLUP and EU predictor
      Generate area effects and bootstrap errors for sampled units as
      $u_{d}^{*(b)}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,\hat{\sigma}_{u,U}^{2}),\quad e_{di}^{*(b)}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,\hat{\sigma}_{e,U}^{2}),\quad i\in s_{d},\ d=1,\ldots,D.$
@@
      With the generated area effects $u_{d}^{*(b)}$ from (i) and the fitted model parameters obtained in step 1, obtain the true bootstrap area means as
      $\mu_{d}^{*(b)}=\bar{\boldsymbol{X}}_{d}^{\prime}\hat{\boldsymbol{\beta}}_{U}+u_{d}^{*(b)},\quad d=1,\ldots,D.$
@@ 7 MSE of EBLUP under FH model
      Taking as true error variances $\hat{\psi}_{d}=\hat{\sigma}_{e}^{2}N_{d}^{-2}\sum_{i\in s_{d}}(w_{di}^{C})^{2}$, $d=1,\ldots,D$, generate, independently, bootstrap area effects and sampling errors as
      $u_{d}^{*(b)}\stackrel{{\scriptstyle iid}}{{\sim}}N(0,\hat{\sigma}_{u}^{2}),\quad e_{d}^{*(b)}\stackrel{{\scriptstyle ind}}{{\sim}}N(0,\hat{\psi}_{d}),\quad d=1,\ldots,D.$
@@
      With $u_{d}^{*(b)}$ generated in (i), obtain the true bootstrap area means as
      $\mu_{d}^{*(b)}=\bar{\boldsymbol{X}}_{d}^{\prime}\hat{\boldsymbol{\beta}}+u_{d}^{*(b)},\quad d=1,\ldots,D.$
@@
      Compute the bootstrap direct estimators
      $\hat{\mu}_{d}^{DIR*(b)}=\mu_{d}^{*(b)}+e_{d}^{*(b)},d=1,\ldots,D.$
@@
We now propose an alternative MSE estimator based on correcting through a bootstrap procedure the Prasad-Rao analytical MSE estimator $\mbox{mse}_{PR}(\hat{\mu}_{d}^{FHD})$ given in (9), by incorporating the uncertainty due to replacing $\psi_{d}$ with $\psi_{d0}=\widehat{\mbox{Var}}_{\pi}(\bar{y}_{dw}^{C}|\mu_{d})$. The bootstrap procedure is the same as the previous one, but additionally, in step 2(iv), calculates the EBLUP based on the “true” bootstrap error variances, $\hat{\boldsymbol{\psi}}=(\hat{\psi}_{1},\ldots,\hat{\psi}_{D})^{\prime}$, $d=1,\ldots,D$, denoted as $\hat{\mu}_{d}^{FHT*(b)}=\hat{\mu}_{d}^{FH*(b)}(\hat{\boldsymbol{\psi}})$. Then, in step 3, it additionally calculates the parametric bootstrap estimator of the MSE of the EBLUP under FH model based on the true error variances, $\hat{\mu}_{d}^{FHT}=\hat{\mu}_{d}^{FH}(\boldsymbol{\psi})$, as follows
@@ 8 Simulation experiments
[…]
@@
[…]
@@
[…]
@@
[…]
@@ 9 Application
First of all, for each area $d=1,\ldots,33$, we calibrate the survey weights $w_{di}$, $i\in s_{d}$, to the known department count $N_{d}$ and the total $X_{1,d}=\sum_{i=1}^{N_{d}}x_{1,di}$, using linear calibration. Now, since both EU predictors assume that BHF model holds for the Maths scores $y_{di}$, we check the model assumptions. Figure 9 in the Appendix shows a histogram and a normal Q-Q plot of unit-level residuals obtained from a REML fit of the model, given by $\hat{e}_{di}=y_{di}-\boldsymbol{x}_{di}^{\prime}\hat{\boldsymbol{\beta}}-\hat{u}_{d}$, for $\hat{u}_{d}=\hat{\gamma}_{d}(\bar{y}_{d}-\bar{\boldsymbol{x}}_{d}^{\prime}\hat{\boldsymbol{\beta}})$ with $\hat{\gamma}_{d}=\hat{\sigma}_{u}^{2}/(\hat{\sigma}_{u}^{2}+\hat{\sigma}_{e}^{2}/n_{d})$. This figure indicates that the normal distribution is tenable, showing only a mild deviation in the tails.

### v2

@@ 2 Area-level FH model
Replacing a consistent estimator $\hat{\sigma}_{u}^{2}(\boldsymbol{\psi})$ of $\sigma_{u}^{2}$ as $D\to\infty$ in (6), we obtain the empirical BLUP (EBLUP) of $\mu_{d}$. Denoting $\hat{\boldsymbol{\beta}}_{FH}(\boldsymbol{\psi})=\tilde{\boldsymbol{\beta}}_{FH}\left(\hat{\sigma}_{u}^{2}(\boldsymbol{\psi}),\boldsymbol{\psi}\right)$ and $\hat{\gamma}_{d}(\boldsymbol{\psi})=\gamma_{d}(\hat{\sigma}_{u}^{2}(\boldsymbol{\psi}),\boldsymbol{\psi})$, the EBLUP of $\mu_{d}$ based on the FH model (3) with error variances $\psi_{d}$, $d=1,\ldots,D$, reads
@@ 6 MSE of PEBLUP and EU predictor
      Generate area effects and bootstrap errors for sampled units as u_d^*(b)∼iid N(0,^σ_u,U^2), e_di^*(b)∼iid N(0, ^σ_e,U^2), i ∈s_d, d=1,…,D.
@@
      With the generated area effects $u_{d}^{*(b)}$ from (i) and the fitted model parameters obtained in step 1, obtain the true bootstrap area means as μ_d^*(b)= ¯X_d’^β_U + u_d^*(b), d=1,…,D.
@@ 7 MSE of EBLUP under FH model
      Taking as true error variances $\hat{\psi}_{d}=\hat{\sigma}_{e}^{2}N_{d}^{-2}\sum_{i\in s_{d}}(w_{di}^{C})^{2}$, $d=1,\ldots,D$, generate, independently, bootstrap area effects and sampling errors as u_d^*(b)∼iid N(0, ^σ_u^2), e_d^*(b)∼ind N(0, ^ψ_d), d=1,…,D.
@@
      With $u_{d}^{*(b)}$ generated in (i), obtain the true bootstrap area means as μ_d^*(b)= ¯X_d’^β + u_d^*(b), d=1,…,D.
@@
      Compute the bootstrap direct estimators ^μ_d^DIR*(b)=μ_d^*(b) + e_d^*(b), d=1,…,D.
@@
[…]
@@ 8 Simulation experiments
In this simulation experiment, we consider a population of N=250  [^0] N=250\,000 units, distributed into $D=25$ areas, with $N_{d}=10\,000$ units in each area $d=1\ldots,D$. We consider two continuous auxiliary variables, $x_{q}$, $q=1,2$, whose values are generated as $x_{q,di}\sim\mbox{Gamma}(k_{qd},1)$ with $k_{1d}=5+3d/D$, $k_{2d}=2$, $i=1\ldots,N_{d}$, $d=1\ldots,D$. The vector of true regression coefficients is taken as $\boldsymbol{\beta}=(4,0\mbox{.}5,-0\mbox{.}4)^{\prime}$ and the variances of the area effects and errors are taken, respectively, as $\sigma_{u}^{2}=0\mbox{.}1^{2}$ and $\sigma_{e}^{2}=0\mbox{.}3^{2}$.
@@
The simulation experiment proceeds as follows. For each MC simulation $\ell=1,\ldots,L$, with L=1  [^0] L=1\,000, we generate a population of $y_{di}^{(\ell)}$ values, $i=1,\ldots,N_{d}$, $d=1,\ldots,D$, from the unit-level BHF model in (10) with the above sizes, auxiliary variables and model parameters. From the generated population values, we calculate the true area means $\mu_{d}^{(\ell)}=\sum_{i=1}^{N_{d}}y_{di}^{(\ell)}$, $d=1,\ldots,D$. Then, taking the values $y_{di}^{(\ell)}$ for the sample units, $i\in s_{d}$, $d=1,\ldots,D$, we compute the four estimators of interest, namely $\hat{\mu}_{d}^{DIR(\ell)}$, $\hat{\mu}_{d}^{FHD(\ell)}$, $\hat{\mu}_{d}^{UA(\ell)}$ and $\hat{\mu}_{d}^{U(\ell)}$. The sample units, $s$, the auxiliary information $\boldsymbol{x}_{di}=(1,x_{1,di},x_{2,di})^{\prime}$, $i=1,\ldots,N_{d}$, $d=1,\ldots,D$, as well as the calibrated weights, are held fixed across MC simulations.
@@
The true MSEs of the EU predictors were previously approximated as described above, but now with L=100  [^0] L=100\,000 MC replicates. Then, a new simulation is performed with L=1  [^0] L=1\,000 MC replicates, calculating in each replicate the analytical MSE estimator $\mbox{mse}_{YR}(\hat{\mu}_{d}^{U})$ and the parametric bootstrap estimator $\mbox{mse}_{PB}(\hat{\mu}_{d}^{U})$ given in (28) with $B=500$ bootstrap replicates.
@@
Next we study the alternative estimators of the MSE of FHD. The true MSE of FHD was again previously approximated with L=100  [^0] L=100\,000 replicates. Figure 4 shows the true MSEs, the means across MC simulations of the Prasad-Rao analytical MSE estimator $\mbox{mse}_{PR}(\hat{\mu}_{d}^{FHD})$ given in (9) (labelled PR mse) and of the two alternative parametric bootstrap MSE estimators $\mbox{mse}_{PB1}(\hat{\mu}_{d}^{FHD})$ and $\mbox{mse}_{PB2}(\hat{\mu}_{d}^{FHD})$ given in (29) and (31) (labelled PB1 mse and PB2 mse respectively). This figure shows that PR analytical MSE estimator, which was not designed to account for the uncertainty due to the estimation of the error variances, underestimates seriously the true MSE of FHD estimator for the areas with $n_{d}\leq 10$, while the two proposed bootstrap estimators perform practically without error. For the areas with sample sizes greater than 10, the three MSE estimators perform similarly, becoming nearly equal for large $n_{d}$. Actually, PR mse is the analytical version of the parametric bootstrap MSE estimator of FHT, $\mbox{mse}_{PB}(\hat{\mu}_{d}^{FHT})$ given in (30), which assumes that the error variances $\psi_{d}$, $d=1,\ldots,D$ are known. None of them incorporate the error that arises from the estimation of these variances.
@@ 9 Application
Colombia is divided into $D=33$ administrative and political divisions, 32 departments and one capital district, hereafter called departments for simplicity. The department population sizes $N_{d}$ (number of students taking the test in that department) range from 268 to 87 556. For this application, we select a SRSWOR of 1.5% from each department. The total sample size is n=8 250n=8\,250, where $n_{d}$ varies between 4 and 1 313.